OpenStudy (anonymous):

Find the distance between points A(-5,7) and B(6,-11) to the nearest tenth.

OpenStudy (anonymous):

hi welcome to OS! do you know the distance formula for two points?

OpenStudy (anonymous):

\( d=\sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2} \) where \(A(-5,7)=(x_1, y_1)\ and\ B(6,-11)=(x_2, y_2)\) plug these points in the formula and you'll get it :)

OpenStudy (anonymous):

hi @Rwelch would you like for me to walk you through??

OpenStudy (anonymous):

Im plugging the points into the formula at the moment

OpenStudy (anonymous):

alright!! :)))))

OpenStudy (anonymous):

just let me know if you need help!

OpenStudy (anonymous):

I think I need the help :/ my answer doesn't seem right

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

d=21.095

OpenStudy (anonymous):

\( \sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}\\ \sqrt{(-5 - 6)^2 + (7 +11)^2}\\\sqrt{(-11)^2+(18)^2}\) it seems like we got the same answer :) BUT the question says "to the nearest tenth" so we have to round it

OpenStudy (anonymous):

21.1? @Data_LG2

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

Awesome :) Thanks for helping me

OpenStudy (anonymous):

no problem :)